40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 8 18:35:09 2019, gautam, Update, LSC, Locking prep 
    Reply  Wed Oct 9 18:15:26 2019, gautam, Update, LSC, PRMI 3f locking PRMI_1f.png
       Reply  Wed Oct 9 22:02:58 2019, gautam, Update, LSC, REFL55 whitening issue 
          Reply  Thu Oct 10 01:12:56 2019, gautam, Update, LSC, Locking studies PRFPMIstudies.png
             Reply  Wed Oct 16 11:42:17 2019, gautam, Update, LSC, Poor separation of PRCL/MICH in 3f signals sensMat.pdf
                Reply  Thu Oct 17 11:19:28 2019, gautam, Update, LSC, Locking activity last night MICH2PRCLnulling.pdfconsolidatedSensingMatrices.pdf.zip
                   Reply  Tue Oct 22 00:52:27 2019, gautam, Update, LSC, Locking updates PRMI3f_ALS_Oct21sensMat.pdf
                      Reply  Tue Oct 29 01:24:45 2019, gautam, Update, LSC, More locking updates POP22_feature.pngPRMI3f_ALS_Oct21sensMat.pdf
                         Reply  Tue Oct 29 15:13:19 2019, gautam, Update, LSC, More locking updates PRMI_lockloss.png
                            Reply  Wed Oct 30 01:27:00 2019, gautam, Update, LSC, More locking updates POP_ASS.png
                               Reply  Wed Oct 30 11:53:41 2019, gautam, Update, LSC, MICH loop shape tuning MICH_ELP.pdfMICH_ELP_TFs.pdf
                               Reply  Wed Oct 30 17:08:40 2019, gautam, Update, LSC, POP22 investigation POP22anomaly.pdf
                                  Reply  Mon Nov 4 15:29:47 2019, gautam, Update, LSC, POP signal path POPchain.pdf
                                     Reply  Mon Nov 4 16:06:58 2019, gautam, Update, LSC, POP optical path IMG_8027.JPGlensRealignment.jpgPOPrealigned.pngPOPdither.pngPRMfixed.png
                                        Reply  Tue Nov 5 11:52:27 2019, gautam, Update, LSC, Locking notes PRMI3f_ALS_Nov4sensMat.pdf
                                           Reply  Wed Nov 6 02:08:48 2019, gautam, Update, LSC, Locking updates 
                                              Reply  Wed Nov 6 17:05:45 2019, gautam, Update, LSC, CARM calibration CARMscan.pdfCARMcalib.pdfscan.hdf5.zip
                                              Reply  Wed Nov 6 17:45:34 2019, gautam, Update, LSC, ~ 
                                                 Reply  Mon Nov 18 21:04:38 2019, gautam, Update, LSC, Locking - some ideas 
                                                    Reply  Wed Nov 20 01:07:18 2019, gautam, Update, LSC, Locking - progress PRFPMI.png
                                                       Reply  Wed Nov 20 12:14:17 2019, gautam, Update, LSC, Locking - progress TRX_TRY_comparison.pdf
                                                          Reply  Wed Nov 20 17:20:24 2019, Yehonathan, Update, LSC, QPD Investigation 
                                                             Reply  Wed Nov 20 17:52:00 2019, gautam, Update, LSC, QPD MEDM screen update newLookQPD.pngTRX_TRY_comparison.pdfIMG_8186.JPG
                                                                Reply  Wed Nov 20 21:29:28 2019, gautam, Update, LSC, PRG ~13 PRG13.pdf
                                                                   Reply  Fri Nov 22 00:54:14 2019, gautam, Update, LSC, locking notes ALSnoiseIncrease.pdf
                                                                      Reply  Mon Nov 25 19:11:22 2019, gautam, Update, LSC, ALS noise re-look ALSnoise.pdf
                                                                         Reply  Wed Nov 27 23:24:01 2019, gautam, Update, LSC, No shaking but no inspiration either 
                                                                            Reply  Sun Dec 1 13:38:48 2019, rana, Update, LSC, No shaking but no inspiration either 
                                                                               Reply  Tue Dec 3 00:03:57 2019, gautam, Update, LSC, Green ALS also shows elevated noise with high arm buildup ALSnoiseIncrease_greenBeat.pdf
                                                                                  Reply  Wed Dec 4 19:54:27 2019, gautam, Update, LSC, A look to the past DRFPMI_2016March.pdf
                                                                                     Reply  Thu Dec 5 01:54:39 2019, gautam, Update, LSC, Partial CM board path engaged AOengaged.pdf
                                                                                        Reply  Fri Dec 6 00:02:48 2019, gautam, Update, LSC, What is the correct way to set the 3f offsets? 
                                     Reply  Fri Nov 15 12:08:04 2019, gautam, Update, LSC, POPDC whitening board 
                         Reply  Tue Oct 29 17:40:48 2019, gautam, Update, LSC, More locking updates broadSweep.pdfzoomSweep.pdf
Message ID: 15045     Entry time: Fri Nov 22 00:54:14 2019     In reply to: 15041     Reply to this: 15046
Author: gautam 
Type: Update 
Category: LSC 
Subject: locking notes 

[KA, GV]

There was no shaking (that disturbed the locking) tonight!

  1. REFL165 Demod phase was adjusted from -111deg to -125deg. To minimize coherence b/w MICH and PRCL.
  2. MICH 3f loop gain changed to 0.3.
  3. If the POP mode shape looks weird, it probably means that the PRM is sligntly misaligned. Tweaking the alignment improves PRMI stability and also makes the arm buildup higher.
  4. Ditto for MICH - slightly touching up the BS alignment can lower ASDC.
  5. Main finding tonight was that the ALS noise seems to get degraded as a function of the CARM offset! As a result of this, CARM goes through several linewidths, and the arm transmission fluctuates wildly.
    • We suspect some scattered light shenanigans. It is not clear to me why this is happening. Possibilities:
    • Scattered ETM transmission somehow makes it into the fiber coupler and degrades the ALS noise.
    • Sacttered ETM transmission makes it onto the Green PDH photodiode and degrades the ALS noise.
    • Backscatter into the PSL degrades the ALS noise.
    • Shadow sensors of either the ITMs, ETMs, BS, or PRM don't have 1064nm filters and get scatterd light, making the cavity length noise worse.
    • Other possibilities?

The problem is hard to debug because we are feeding back on the ETMs, BS and PRM, and at the low CARM offset (= high PRG), all the DoFs are cross coupled strongly so just by looking at error/control signals, I can't directly determine where the noise is originating. The fact that the ALS CARM spectrum shows a noise increase suggests that the problem has to do with the test masses, PSL, IMC, or end green PDH setups.

My plan is to do a systematic campaign and eliminate some of these possibilities - e.g. install some baffling around the fiber coupler and the end green PDH photodiodes and see if there is any improvement in the situation.

* In attachment #1, the "Ref" traces are when the CARM offset is 0, and the arms are buzzing in and out of resonance. The non-reference traces are for when the CARM offset is ~28kHz (so several linewidths away from resonance).

Attachment 1: ALSnoiseIncrease.pdf  108 kB  Uploaded Fri Nov 22 16:41:58 2019  | Hide | Hide all
ALSnoiseIncrease.pdf
ELOG V3.1.3-